OpenStudy (tainted):

Which of the following is the vertex of the quadratic equation y = 6x2 – 24x + 32?

OpenStudy (anonymous):

y-8=6(x^2-4x+4) or (y-8)=6(x-2)^2 hence vertex of the above parabola is at (2,8)

OpenStudy (anonymous):

use vertex formula \[\frac{ -b }{ 2a }\]

OpenStudy (anonymous):

b=-24 and a=6

OpenStudy (anonymous):

\[\frac{ -(-24) }{ 2(6) }=2\]

OpenStudy (anonymous):

then: \[6(2)^{2}-24(2)+32=8\]

OpenStudy (anonymous):

(2,8)